OpenStudy (anonymous):

let p, q be natural numbers and P(p,q) be some statement true for some pairs p,q but not for others does this being true: ∃p∀q:P(p,q) imply that the following statement is true: ∀q∃p:P(p,q)

OpenStudy (anonymous):

can you be my @Hero and help me?

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@AccessDenied

OpenStudy (anonymous):

im sorry for the tagging spree

OpenStudy (anonymous):

@bahrom7893

OpenStudy (anonymous):

Let me make sure I understand the statement: Does "There Exists some p for All q such that P(p,q) is true" imlpy that "For All q, There Exists some p such that P(p,q) is true"

OpenStudy (anonymous):

surely it does

OpenStudy (anonymous):

Because it sure seems like it to me.

OpenStudy (anonymous):

on my worksheet it asks for a counter example of this statement and i haven't found one. also i believe a proof is available

OpenStudy (anonymous):

If I translated it correctly, I can't see a counterexample.

OpenStudy (anonymous):

Nevermind you're right.

OpenStudy (anonymous):

There Exists some p for All q such that P(p,q) is true let this value of p be denoted as p* now For All q, There Exists some p* such that P(p,q) is true

OpenStudy (anonymous):

im confused though, my worksheet asks effectively for a counterexample

OpenStudy (anonymous):

The latter does not imply the former, however, so make sure you read it properly.

OpenStudy (anonymous):

Is the empty set a counter example?

OpenStudy (anonymous):

Or is the counterexample the empty set?

OpenStudy (anonymous):

here is the original question (its part (b) ) Let P(p; q) denote a statement about natural numbers p and q which is true for some pairs p and q and false for other pairs. For example, P(p; q) might be the statement `p < q^2 ' or the statement `q = 10^(23p) + 42'. Give examples in which (a) ∀p∃q P(p; q) is true but ∃q∀p P(p; q) is false; (b) ∃p∀q P(p; q) is true but ∀q∃p P(p; q) is false. \[\exists \forall\]

OpenStudy (anonymous):

ignore the two symbols at the bottom

OpenStudy (anonymous):

The first is incorrect... for all p there exists a q such that P(p;q) does not imply that there exists q such that for all p P(p;q). The second looks fine to me though.

OpenStudy (anonymous):

yeah an example for the first is pq = q^2